### v1

# Meaningful Learning: Advancing Abstract Reasoning in Large Language Models via Generic Fact Guidance

Abstract
Large language models (LLMs) have developed impressive performance and strong explainability across various reasoning scenarios, marking a significant stride towards mimicking human-like intelligence. Despite this, when tasked with simple questions supported by a generic fact, LLMs often fail to provide consistent and precise answers, indicating a deficiency in abstract reasoning abilities. This has sparked a vigorous debate about whether LLMs are genuinely reasoning or merely memorizing. In light of this, we design a preliminary study to quantify and delve into the abstract reasoning abilities of existing LLMs. Our findings reveal a substantial discrepancy between their general reasoning and abstract reasoning performances. To relieve this problem, we tailor an abstract reasoning dataset (AbsR) together with a meaningful learning paradigm to teach LLMs how to leverage generic facts for reasoning purposes. The results show that our approach not only boosts the general reasoning performance of LLMs but also makes considerable strides towards their capacity for abstract reasoning, moving beyond simple memorization or imitation to a more nuanced understanding and application of generic facts.
1 Introduction
Humans proceed meaningful learning to induce common patterns or high-level abstractions to acquire abstract reasoning abilities Lake et al. (2017); Qiu et al. (2023). Such capabilities allow us to apply broad principles across diverse situations, demonstrating a deep understanding and versatile application of knowledge. As shown in Figure 1 (a), humans can deduce “rock dissolved” and “solution turns orange” when given “adding rock into hydrochloric acid” and “adding rust into sulphuric acid”, respectively. This stems from the established expertise in human mind (the generic fact “acid is corrosive”), and extends to applications in different scenarios. Of late, remarkable headways of LLMs have pushed AI much further towards human-like intelligence Achiam et al. (2023); Touvron et al. (2023). Interestingly, can LLMs act like humans to instinctively consider the generic fact for flexible applications in different scenarios?
Figure 1: The responses of (a) Humans and (b) LLMs when facing two questions supported by a genric fact.
LLMs Size Vanilla Accuracy AbsAcc
LLaMA-2 7B 51.67 24.82
13B 68.45 42.94
Orca-2 7B 74.23 51.19
13B 79.83 59.47
Table 1: The vanilla accuracy and abstract reasoning accuracy (AbsAcc) of LLaMA-2 Touvron et al. (2023) and Orca-2 Mitra et al. (2023) on e-CARE Du et al. (2022). We will formally define AbsAcc in Sec 2.1.
In our pilot investigation, we observed that LLMs appear to lack satisfying capabilities in abstract reasoning. As illustrated in Table 1, there is a notable discrepancy—exceeding 20%—between vanilla accuracy and abstract reasoning accuracy (AbsAcc). Despite the extensive pre-training of LLMs, which equips them with a vast repository of generic facts, they seem unable to leverage this information as flexibly as humans do (Figure 1 (b)).
In this paper, we aim to systematically and quantitatively investigate the abstract reasoning of LLMs. We first formally define the abstract reasoning metric. Then we carry out abstract reasoning and knowledge probing tasks for a preliminary study. Our investigation comprises two main components: the reasoning tasks can quantify the abstract reasoning capabilities of LLMs. The generic fact probing task can give a further analysis from the perspective of generic fact mastery. Hereafter, we improve abstract reasoning of LLMs based on the analysis of the preliminary study. Unlike existing chain-of-thought (CoT) methods Wei et al. (2022a); Kojima et al. (2022), which conduct reasoning with uncontrolled step-by-step explanation, we create a dataset AbsR tailored for abstract reasoning. AbsR not only includes generic facts but also their guided explanations, offering a coached approach to understand the process of reasoning with common patterns. Finally, to make LLMs subconsciously exploit the generic fact like humans, we design a simple but effective learning paradigm called meaningful learning to simulate the implicit knowledge learning process Du et al. (2019), enabling LLMs to implicitly learn and utilize generic facts for reasoning under the guidance of generic facts and the generic-fact-guided explanation.
Extensive experiments across multiple reasoning and natural language understanding benchmarks, demonstrate that our proposed paradigm for meaningful learning not only improves general reasoning of LLMs, but also excels in abstract reasoning. This distinction in performance is particularly pronounced in the realm of abstract reasoning, underscoring the unique efficacy of our approach in nurturing the higher-order thinking skills that are essential for sophisticated cognitive processing in artificial intelligence systems. Further analysis and ablation studies yield additional evidence supporting the effectiveness of our approach.
We summarize our contributions as follows:
  • We provide a systematic and quantitative analysis of abstract reasoning in current LLMs, and we develop an abstract reasoning dataset with generic-fact-guided explanation.
  • We propose a meaningful learning paradigm, which can stimulate the abstract reasoning capabilities of existing LLMs.
  • We achieve significant improvement in general reasoning and abstract reasoning on a wide range of benchmarks.
Figure 2: The computation of abstract reasoning metric.
2 Abstract Reasoning Study
In this section, we initially define the evaluation metric of abstract reasoning to establish quantifiable standards. Then we introduce the LLMs used in this study. Finally, we conduct two experiments: one to directly estimate the abstract reasoning abilities of LLMs, and the other to further analyze it from the perspective of general fact mastery.
2.1 Abstract Reasoning Metric
Abstract reasoning requires models to apply general patterns of high-level abstractions to different scenarios or questions. Different from general reasoning, which focuses on each single example, abstract reasoning takes the examples supported by a generic fact as a whole. We define abstract reasoning accuracy due to the lack of a proper metric.
Given a dataset $D$ with $n$ generic facts $R=\{r_{1},r_{2}\cdots,r_{n}\}$, for the $i$-th generic fact $r_{i}$, it supports $m_{i}$ examples $S_{i}=\{s^{i}_{1},s^{i}_{2},\cdots,s^{i}_{m_{i}}\}\in D$.
As shown in Figure 2, for hard constraint, akin to Qiu et al. (2023), we suppose an LLM $\mathcal{LM}$ has grasped the generic fact $r_{i}$ if and only if $\mathcal{LM}$ can correctly answer all examples in $S_{i}$. Hence, we define the abstract reasoning accuracy as the proportion of generic facts that $\mathcal{LM}$ has grasped:
$\texttt{AbsAcc}=\frac{\sum_{i}\mathbb{I}(\Phi(L,S_{i})=|S_{i}|)}{n},$ (1)
where $\Phi$ quantifies the number of examples in $S_{i}$ that $\mathcal{LM}$ answers correctly. $\mathbb{I}$ serves as the indicator function, taking 1 when the specified condition is met and 0 otherwise.
2.2 Large Language Models
We offer an overview of the LLMs employed in our preliminary and upcoming experiments:
[…]
2.3 Experiment I: Abstract Reasoning
We formulate the first preliminary experiment to gain an initial insight into the abstract reasoning abilities of current LLMs.
Examples Generic Facts
Original Filtered Orginal Filtered
21,314 13,877 13,045 5,608
Table 2: The statistics of original and filtered e-CARE.
To achieve this goal, we choose a multiple-choice question answering dataset e-CARE Du et al. (2022) for implementation. e-CARE is a large-scale explainable causal reasoning dataset with a generic fact in each example, and a generic fact can support more than one example. The full set (train, test, and dev) of e-CARE is selected for experiments. We filter out the examples that the corresponding generic facts support only one example. Table 2 shows the statistics of e-CARE.
We follow zero-shot setting to prompt the LLMs for answer selection from multiple choices, and the generic facts are not provided in the inputs (refer to Appendix A.1). Table 1 shows the overall results:
(1) There is a significant disparity (more than 20%) between the vanilla accuracy and AbsAcc across all LLMs. This proves two aspects: on the one hand, vanilla accuracy is not ideal for estimating abstract reasoning. On the other hand, although LLMs could answer questions accurately, they still cannot capture the common patterns or generic facts behind the questions to reach the answers.
(2) Larger LLMs outperform smaller ones due to their richer and more accurate knowledge. Orca-2, fine-tuned on LLaMA-2 with extensive reasoning examples, notably outperforms LLaMA-2, indicating that supervised fine-tuning enhances both reasoning and abstract reasoning abilities. However, even with large-scale and reasoning-specific post-training, the significant gap between vanilla accuracy and AbsAcc remains unbridged.
2.4 Experiment II: Generic Fact Probing
To delve into the abstract reasoning from the perspective of generic fact matery, we design a generic fact probing task to ask LLMs whether they possess the given generic fact (refer to Appendix A.2). Drawing inspiration from knowledge uncertainty estimation Amayuelas et al. (2023), we formulate this task as a yes or no problem.
LLMs LLaMA-2 Orca-2
7B 13B 7B 13B
Accuracy 62.44 98.86 26.97 12.78
Table 3: The overall results of the generic facts probing.
We also use e-CARE and the above LLMs to carry out experiments. The overall results are shown in Table 3. We could conclude as follows:
(1) Both 7B and 13B LLaMA-2 have substantial knowledge reserves, with LLaMA-2-13B being particularly knowing nearly all generic facts in e-CARE. This indicates the poor AbsAcc might come from a lack of ability to utilize knowledge.
(2) In contrast, Orca-2 knows limited generic facts, and the larger the model, the less it knows. This contrasts sharply with its strong reasoning abilities (vanilla accuracy). We suppose training Orca-2 is more like rote learning rather than meaningful learning, which implies Orca-2 tends to lose knowledge while acquiring reasoning.
Category Metrics LLaMA-2 Orca-2
7B 13B 7B 13B
Known AbsAcc 34.26 43.51 64.98 81.51
Unknown AbsAcc 17.19 41.84 49.00 58.38
Table 4: The categorized abstract reasoning accuracy based on whether the generic facts are known.
To delve into the relation between generic facts and abstract reasoning, we categorize the examples in e-CARE according to whether LLMs know the generic facts. The results are represented in Table 4, from which we can draw the following conclusions:
(1) The performance of “Known” category has superiority over “Unknown” category across all LLMs. This demonstrates knowing the generic facts is helpful for abstract reasoning, which provides a potential avenue for abstract reasoning enhancement through generic fact retrieval.
(2) Nonetheless, most of the results in “Unknown” category still have undeniable performance, suggesting that LLMs may engage in reasoning based on suspicious correlations or rely on memory recall. We can reasonably speculate that LLMs may consciously conduct memorization when answering reasoning-related questions.
In a conclusion, we can enhance the abstract reasoning abilities of LLMs in two ways: (1) Knowledge, injecting generic facts into the parametric memory of LLMs or offering LLMs the generic fact through prompts; (2) Reasoning, teaching LLMs how to utilize the generic facts to provide more precise responses to questions.
Figure 3: Two examples of AbsR guided by the generic fact about “more powerful computer”.
3 AbsR: Abstract Reasoning Dataset
To enhance the abstract reasoning abilities of LLMs through Knowledge and Reasoning ways, we need a dataset that can not only provide generic facts, but also teach LLMs how to exploit the generic fact to derive the correct answer in different scenarios (meaningful learning). Since there is no suitable dataset to assist us in achieving the goals, we create an Abstract Reasoning dataset (AbsR).
3.1 Generic Fact Collection
To obtain such a dataset described above, we need to collect the generic fact first. We choose GenericsKB Bhakthavatsalam et al. (2020) as the foundation of AbsR to supply generic facts. GenericsKB is a large-scale knowledge base with over 3.4 million sentences (e.g. Dogs bark) expressing general truths, each of which also includes the corresponding concept (e.g. Dog) and the confidence score (between 0 and 1). Finally, we sample a collection of 4,613 generic facts from GenericsKB (details can refer to Appendix B).
3.2 Dataset Construction
To construct the whole dataset, and inspired by previous works Wang et al. (2022a); Xu et al. (2023), we choose GPT-4 Achiam et al. (2023) as our data annotator. The API we used is gpt-4-1106-preview.
Specifically, for each sampled generic fact $r_{i}$, we would ask GPT-4 to create samples $S_{i}=\{s_{1}^{i},\cdots,s_{m_{i}}^{i}|1\leq m_{i}\leq 3\}$ according to $r_{i}$. Each sample $s^{i}_{j}$ contains the following terms: (1) a question $X^{i}_{j}$ with a few options, (2) a response $Y^{i}_{j}$ with an answer and an explanation guided by $r_{i}$. All terms form a triple $s^{i}_{j}=<r_{i},X^{i}_{j},Y^{i}_{j}>$ for each sample. The prompt we used for dataset creation can refer to Appendix C. An illustration of the created examples is shown in Figure 3.
Finally, without loss of generality, given the $j$-th sample $s^{i}_{j}$ of generic fact $r_{i}$, we can create two kinds of examples for our proposed meaningful learning. One is predicting $Y^{i}_{j}$ given $<r_{i},X^{i}_{j}>$ (K-example), while the other is predicting $Y^{i}_{j}$ given only $X^{i}_{j}$ (R-example). The examples are used to implicitly enhance the abstract reasoning abilities of LLMs through Knowledge and Reasoning ways. Table 5 shows the statistics of AbsR.
Examples Questions Generic Facts
Train 18,020 14,010 4,613
Test 200 200 104
Table 5: The statistics of AbsR.
4 Method: Meaningful Learning
To imitate humans’ instinctive use of general facts in reasoning, we develop a simple but effective learning paradigm called meaningful learning (MeanLearn). It can enhance the abstract reasoning abilities of LLMs in Knowledge and Reasoning ways. This is inspired by the implicit knowledge learning process, which uses hidden variables to learn event background knowledge Du et al. (2019). MeanLearn can make LLMs implicitly learn generic facts and solve problems under the guidance of generic facts.
Specifically, given an LLM $\mathcal{LM}$ with parameters $\theta$ and the input $x$, which constructs a conditional probability for the outputs $y$:
$\mathcal{LM}(x,y,\theta)=-\sum_{t}\log p_{\theta}(y_{t}|x,y_{<t}).$ (2)
[…]
$\mathcal{LM}(X,Y,\theta)=-\sum_{t}\log p_{\theta}(Y_{t}|X,Y_{<t}),$ (3)
$\mathcal{LM}(X,r,Y,\theta)=-\sum_{t}\log q_{\theta}(Y_{t}|X,r,Y_{<t}).$
Hereafter, on the one hand, we train $\mathcal{LM}$ to learn policies $p_{\theta}(Y_{t}|X,Y_{<t})$ and $q_{\theta}(Y_{t}|X,r,Y_{<t})$ together to push them closer, enabling $\mathcal{LM}$ implicitly learn the generac fact $r$. On the other hand, $\mathcal{LM}$ can learn how to apply $r$ by learning to generate the explanation in $Y$. Finally, we can obtain better abstract reasoning abilities when only given the question $X$, just like humans instinctively answer questions with general facts.
5 Experiments
5.1 Training Details
We use LoRA Hu et al. (2022) for parameter effcient finetuning of Orca-2. The batch size is set as 256 and 240 for 7B and 13B Orca-2, respectively. We train them for one epoch with a learning rate of 5e-5. 7B and 13B Orca-2 are respectively trained on 2 and 3 NVIDIA A100 80GB PCIe GPUs.
5.2 Evaluation Benchmarks
To evaluate the MeanLearn paradigm, in addition to AbsR, we also include a wide range of benchmarks of reasoning and natural language understanding:
[…]
5.3 Evalution Settings
Since the generation-based evaluation is time-consuming due to the limitations of computing resources, we follow OpenCompass Contributors (2023) to have hybrid evaluation criteria:
(1) For free-form generation tasks (4 tasks in BBH), we ask LLMs to generate no more than 250 tokens with a temperature of 0.0, and then we utilize exact match for evaluation.
(2) For the other classification and mutipul-choice questions, perplexity (PPL) of each category or option is adopted for prediction. The evaluation templates can refer to Appendix E. Take Orca-2 and MeanLearn as an example, we use:
<|im_start|> system{INSTRUCTION} <|im_end|> <|im_start|> userQuestion:{QUESTION} <|im_end|> <|im_start|> assistantAnswer:{OPTION} <|im_end|>
The text enclosed within curly braces is a placeholder. {INSTRUCTION} will be filled with task-specific instruction. {QUESTION} will be filled with the question and all candidate options or categories. {OPTION} will be filled with the option index only (e.g. A).
To measure the performance, the option or category with the lowest PPL will be chosen as the prediction of corresponding LLM. The performance of each benchmark is weighted across all tasks of datasets in the benchmark.
Size LLMs AbsR AGIEval RACE BBH Commonsense MMLU ARC-e ARC-c Average
Vanilla Accuracy
7B LLaMA-2 50.00 27.89 38.85 26.00 55.37 41.65 58.73 43.05 42.69
Orca-2 73.50 38.57 73.32 34.51 71.58 50.11 79.19 73.90 61.84
MeanLearn 77.00 38.33 77.22 35.71 76.55 53.04 86.60 78.31 65.35
13B LLaMA-2 73.50 34.41 59.46 30.61 61.00 51.87 71.25 55.25 54.67
Orca-2 66.50 43.75 66.86 39.39 65.92 47.27 82.19 70.85 60.34
MeanLearn 66.00 43.95 70.47 40.45 69.93 51.62 89.59 82.03 64.26
AbsAcc
7B LLaMA-2 35.42 16.09 14.08 9.56 47.42 25.48 50.27 35.51 29.23
Orca-2 60.42 26.27 50.90 15.67 64.02 33.68 72.16 67.76 48.86
MeanLearn 64.58 25.27 57.10 16.67 71.17 37.24 81.35 73.83 53.40
13B LLaMA-2 61.46 21.01 32.17 13.74 52.87 35.72 62.43 44.86 40.53
Orca-2 50.00 29.50 42.12 20.61 56.41 31.11 75.95 62.15 45.98
MeanLearn 45.83 30.06 47.89 22.00 61.06 35.24 85.41 77.10 50.57
Table 6: The overall vanilla accuracy and AbsAcc (%) of all LLMs.
5.4 Results: Vanilla Accuracy
Table 6 shows the Vanilla Accuracy of MeanLearn and baselines on AbsR and the benchmarks. We can infer that:
(1) Orca-2 siginificantly outperforms LLaMA-2 on most benchmarks, which indicates that massive reasoning-specific data can teach LLMs to develop much better reasoning abilities.
(2) MeanLearn has superiority over LLaMA-2 and Orca-2 on nearly all benchmarks. This demonstrates MeanLearn can use generic facts to effectively guide LLMs to reason more properly and logically, resulting in better reasoning capabilities.
(3) The more challenging the benchmark (e.g. BBH), the less pronounced the improvement in the performance of MeanLearn. This is mainly because the improvement is mostly constrained by the complex scenarios in harder benchmarks, which requires mightier base LLMs or more complex training data to perform and learn multi-step and compositional reasoning.
(4) Post-training (Orca-2 and MeanLearn) does not always bring improvements (e.g. AGIEval). The primary reason is the reasoning-specific post-training, which is bound to affect the general abilities and the handle of uncovered reasoning types.
(5) Larger LLMs do not always yield better results. We suppose the main factors behind this are in two ways: on the one hand, LLMs with different parameter sizes might be good at different tasks, excessive contemplation may lead to decreases on some tasks Wei et al. (2022b). On the other hand, the perplexity-based evaluation setting might has moderate cross-scale generalization ability.
5.5 Results: Abstract Reasoning Accuracy
Since the generic facts are not given in the benchmarks, we train a RoBERTa-Large Liu et al. (2019) with constrastive learning to cluster examples supported by the same generic fact. In particular, we utilize e-CARE as the training set, examples guided by a same generic fact will be pushed closer, or they will be pushed away. Cosine similarity is used to measure the distance between examples. For clustering, examples sharing a similarity above 0.6 are considered to be supported by the same generic fact. Each cluster contains no more than 3 examples.
Hereafter, we calculate the abstract reasoning metrics based on the results of clustering. The overall results are shown in Table 6 ( AbsAcc), we can have the following observations:
(1) The AbsAcc is much lower than the vanilla accuracy, indicating the large gap between general reasoning and abatract reasoning. Perhaps more training data can help to fill the gap between vanilla accuracy and abstract reasoning accuracy.
(2) MeanLearn surpasses Orca-2 and LLaMA-2, exhibiting greater advantages in AbsAcc than in vanilla accuracy. This suggests that the improvement in vanilla accuracy is more likely due to the enhancement of abstract reasoning abilities. MeanLearn can implicitly tell LLMs the generic fact and teach them how to utilize it, encouraging LLMs focus more on abstract reasoning.
6 Analysis
To further investigate the pros, cons, and effectiveness of MeanLearn and baselines, we conduct several in-depth analyses and ablation studies.
Figure 4: Visualization of reasoning capabilities on MMLU, which is categorized by task domians.
Size LLMs AbsR AGIEval RACE BBH Commonsense MMLU ARC-e ARC-c Average
Vanilla Accuracy
7B MeanLearn 77.00 38.33 77.22 35.71 76.55 53.04 86.60 78.31 65.35
w/o Knowledge 73.50 38.89 74.10 33.80 73.28 50.97 79.72 75.25 62.44
w/o Reasoning 72.50 38.37 72.60 33.39 71.68 49.55 77.25 73.22 61.07
w/o Guidance 73.50 38.53 73.14 33.65 73.57 50.96 80.60 75.25 62.40
13B MeanLearn 66.00 43.95 70.47 40.45 69.93 51.62 89.59 82.03 64.26
w/o Knowledge 63.50 44.07 67.19 36.80 70.45 47.88 84.13 74.58 61.08
w/o Reasoning 56.00 43.95 58.05 35.91 68.09 45.31 82.54 69.83 57.46
w/o Guidance 66.50 43.75 66.86 36.92 65.92 47.27 82.19 70.85 60.03
AbsAcc
7B MeanLearn 64.58 25.27 57.10 16.67 71.17 37.24 81.35 73.83 53.40
w/o Knowledge 61.46 26.67 52.91 14.74 66.26 34.84 72.70 69.16 49.84
w/o Reasoning 59.38 26.18 49.66 14.15 64.13 33.27 69.19 67.29 47.91
w/o Guidance 61.46 26.44 52.24 14.76 55.56 34.88 74.05 69.16 48.57
13B MeanLearn 45.83 30.06 47.89 22.00 61.06 35.24 85.41 77.10 50.57
w/o Knowledge 46.88 30.43 42.87 18.83 62.95 32.10 78.11 66.82 47.37
w/o Reasoning 36.46 30.81 31.19 18.26 59.70 29.12 75.68 61.68 42.86
w/o Guidance 50.00 29.58 42.12 18.82 56.41 31.11 75.95 62.15 45.77
Table 7: The overall reasults of ablation studies. “w/o” denotes without.
6.1 Performance on Different Domains
For potential future improvements, we categorize tasks of MMLU based on their respective domains to explore the domains where MeanLearn and baselines excel and where they does not excel. The results are shown in Figure 4 with a total of 16 categories (e.g. law), from which we can conclude:
(1) All LLMs possess poor vanilla accuracy and AbsAcc in math and chemistry. We conjecture the primary factors are the complexity of math problems and the shortage of chemistry knowledge.
(2) Compared to Orca-2, MeanLearn possesses considerable advantages in engineering, geography, and physics & astronomy domains. We suppose this is related to the domain coverage of AbsR.
(3) Since the overall AbsAcc is not relatively high, for potential improvement in the future, apart from focusing more on the domains with poor vanilla accuracy and AbsAcc, we should also increase the overall volume of finetuning data.
6.2 Effect of Knowledge and Reasoning
We utilize Knowledge and Reasoning ways to enhance abstract reasoning in LLMs. To investigate the effect of them, we conduct an ablation study of AbsR to respectively remove the Knowledge (generic facts) and Reasoning (explanations) in each example. Results are shown in Table 7 (w/o Knowledge and w/o Reasoning), we can observe:
(1) Removing Knowledge (generic facts) or Reasoning (explanation) leads to a decrease in performance. This is because the generic facts and their guided reasoning process in AbsR can teach LLMs to learn and utilize the generic fact behind the questions to reason more logically and properly.
(2) In MeanLearn, Knowledge contributes less in boosting the vanilla and abstract reasoning than Reasoning. It indicates the importance of teaching LLMs how to reason under the guidance of generic facts. Furthermore, it reveals the bottleneck of small-scale LLMs reasoning is caused more by reasoning itself rather than knowledge.
(3) MeanLearn without Knowledge can defeat Orca-2 on vanilla accuracy and AbsAcc, while MeanLearn without Reasoning cannot. This implies the synergy effects between Knowledge and Reasoning (MeanLearn) can push LLMs further.
(4) In AGIEval, Knowledge and Reasoning are less critical due to its focus on complex reading comprehension without needing vast knowledge but. Moreover, Knowledge seems useless in Commonsense. When providing LLMs with knowledge, we should also teach them how to use it.
6.3 Effect of Guided Explanations
When constructing AbsR, we would ask GPT-4 to generate an explanation guided by the generic fact (Sec 3.2). Hence, we devise another ablation study to investigate the effect of the guidance brought by the generic fact.
To be precise, we follow the process of generating samples with GPT-4 in Sec 3.2, but we do not provide the generic fact in the prompt. Therefore, the explanation is guided by GPT-4 itself. We denote this dataset as $*{}^{*}$. Results are shown in Table 7 (w/o Guidance), from which we can infer:
(1) Training 7B and 13B Orca-2 with $*{}^{*}$ can still outperform Orca-2 on vanilla accuracy. However, it cannot bring stable improvements on AbsAcc. This underscores that reasoning-specific post-training can boost performance, while improving abstract reasoning performance might depend on the difficulty of the benchmarks.
(2) MeanLearn without Knowledge outperforms MeanLearn without Guidance, showing the superiority of training with generic-fact-guided explanation. This stems from the generic-fact-guided explanation that can teach LLMs to use the regular pattern behind the questions for better reasoning.
7 Related Work
7.1 Reasoning with LLMs
LLMs have overturned the research paradigm of NLP, which makes reasoning more accessible. Using LLMs to perform reasoning can be categorized into tunning-based and prompt-based methods.
[…]
As for prompt-based methods, Wei et al. (2022b) introduced CoT prompt to elicit reasoning in LLMs by encouraging step-by-step thinking. Additionally, Kojima et al. (2022) presented zero-shot CoT to bypass manual annotation issues. Subsequently, Zhang et al. (2023) proposed AutoCoT to automatically generate few-shot examples with CoT for reasoning. Various CoT-based methods have since emerged, such as complex CoT Fu et al. (2023), and tree of thought Yao et al. (2023), etc.
These methods are designed for improving the general reasoning abilities of LLMs, while we focus on the abstract reasoning capabilities of LLMs.
7.2 LLMs-as-Annotators
LLMs are increasingly employed for data labeling due to their advanced ability to follow instructions.
Some studies distilled various CoT data from LLMs to train student models (Li et al., 2023; Ho et al., 2023; inter alia), achieving impressive performance. For instance, Li et al. (2023) distilled symbolic CoT from GPT-3 and observed enhanced performance in commonsense reasoning. With the advent of open-source LLMs like LLaMA Touvron et al. (2023), other studies distilled instruction data from large-scale LLMs to train small-scale LLMs (Xu et al., 2023; Mukherjee et al., 2023; inter alia). For example, Xu et al. (2023) employed an evolving strategy to obtain complex data, training LLaMA and achieving strong performance.
We employ LLMs to generate training data and develop AbsR tailored for abstract reasoning.
8 Conclusion
In this paper, we investigate the abstract reasoning of open-source LLMs. To achieve this, we first define the evaluation metric of abstract reasoning and design a series of preliminary experiments. Then we discover a large performance gap between general reasoning and abstract reasoning. Hence, we tailor an abstract reasoning dataset AbsR with the help of GPT-4 to enhance LLMs. Finally, we devise a simple but effective paradigm (MeanLearn) to teach LLMs abstract reasoning in Knowledge and Reasoning ways. Extensive experiments demonstrate the superiority and effectiveness of MeanLearn in vanilla and abstract reasoning accuracies.
Limitations
There are several limitations of our work: First, the size of the training data should be larger to obtain better and more stable results. Second, since the large performance improvement might be due to the PPL-based evaluation, the evaluation setting should be in a generative way to get more precise results. There should be some examples (whether MeanLearn can generate meaningful explanations) to demonstrate the superiority of MeanLearn. Third, more efforts should be made to focus more on abstract reasoning itself. Finally, more LLMs should be incorporated for training to demonstrate the generalizability of MeanLearn.
Appendix A Asbstract Reasoning Study
A.1 Prompt for Experiment I
<|im_start|> system You are Orca, an AI language model created by Microsoft. You are a cautious assistant. You carefully follow instructions. You are helpful and harmless and you follow ethical guidelines and promote positive behavior.<|im_end|> <|im_start|> user{premise} Hypothesis1 or Hypothesis2? Hypothesis1: {hypothesis1} Hypothesis2: {hypothesis2} Your answer should follow the format like ‘‘Answer: Hypothesis(1 or 2) is more plausible. Explanation: ___’’<|im_end|> <|im_start|> assistant
The placeholders {premise}, {hypothesis1}, and {hypothesis2} will be filled with the corresponding terms in each example of e-CARE dataset.
A.2 Prmpt for Experiment II
<|im_start|> system You are Orca, an AI language model created by Microsoft. You are a cautious assistant. You carefully follow instructions. You are helpful and harmless and you follow ethical guidelines and promote positive behavior.<|im_end|> <|im_start|> user Question: {premise} Hypothesis1 or Hypothesis2? Hypothesis1: {hypothesis1} Hypothesis2: {hypothesis2} Do you think {fact} can be used to answer this question? You answer should follow the format like ‘‘Answer: Yes or No.’’<|im_end|> <|im_start|> assistant
The placeholders {premise}, {hypothesis1}, {hypothesis1}, and {fact} will be filled with the corresponding terms in each example of e-CARE dataset.
Appendix B Details for Generic Facts Filtering
We apply the following steps to filter generics fact from GenericsKB:
(1) Exclude sentences with confidence $<$ 0.7.
(2) Categorize the sentences by the concepts.
(3) Randomly sample 4,613 concepts, and then randomly sample one sentence as the generic fact from the category of each sampled concept.
Hereafter, we can obtain a collection of 4,613 generic facts.
Appendix C Prompt for AbsR Construction
We use the following prompt to generate our AbsR examples.
{number}
The placeholder {number} will be randomly filled with “one”, “two” or “three”.
Appendix D K-example and R-example
D.1 K-example
<|im_start|> system You are a cautious assistant. You carefully follow instructions. You are given a question, a few options, and a rule, you should follow the given rule to answer the question. Your response should follow the format like ‘‘Explanation: ___Answer: ____’’<|im_end|> <|im_start|> user Fact: Cookie files are simple text files that can be viewed in Windows Notepad. Question: What type of application can be used to open and view cookie files on a Windows computer? Options: A) Microsoft Excel B) Adobe Photoshop C) Windows Notepad D) 3D Modeling Software<|im_end|> <|im_start|> assistant Explanation: Windows Notepad is a text editor that is capable of opening and displaying the contents of simple text files, which is the format of cookie files. Answer: C) Windows Notepad<|im_end|>
D.2 R-example
<|im_start|> system You are a cautious assistant. You carefully follow instructions. You are given a question together with a few options, you should give an explanation first then answer the question. Your response should follow the format like ‘‘Explanation: ___Answer: ____’’<|im_end|> <|im_start|> user Question: What type of application can be used to open and view cookie files on a Windows computer? Options: A) Microsoft Excel B) Adobe Photoshop C) Windows Notepad D) 3D Modeling Software<|im_end|> <|im_start|> assistant Explanation: Windows Notepad is a text editor that is capable of opening and displaying the contents of simple text files, which is the format of cookie files. Answer: C) Windows Notepad<|im_end|>
Appendix E Templates for Evaluation
E.1 LLaMA-2
The following are multiple choice questions (with answers) about abstract algebra. Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q. A. 0 B. 4 C. 2 D. 6 Answer: A
E.2 Orca-2 and MeanLearn
<|im_start|> system The following are multiple choice questions (with answers) about abstract algebra.<|im_end|> <|im_start|> user Question: Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q. Options: A. 0 B. 4 C. 2 D. 6<|im_end|> <|im_start|> assistant Answer: A<|im_end|>